OpenStudy (anonymous):

Find the product of (x + 3)(x - 3).

OpenStudy (anonymous):

!!!!!!!!!!!!!!!!!!!medal and fann!!!!!!!!!!!!!!!!!!!!!

OpenStudy (onepieceftw):

Use foil. First term, outer term, inner term, last term.

OpenStudy (onepieceftw):

x times x x times -3 3 times x 3 times -3

OpenStudy (anonymous):

x^2 - 6x + 9 x^2 + 6x + 9 x^2 + 9 x^2 - 9

OpenStudy (onepieceftw):

Very good!

OpenStudy (anonymous):

those r the choices

OpenStudy (onepieceftw):

You can do this calculation. Which one do you think it is?

OpenStudy (anonymous):

not c or d

OpenStudy (anonymous):

and it is negative s a ?

OpenStudy (anonymous):

so*

OpenStudy (onepieceftw):

Try again.

OpenStudy (anonymous):

3*3 is 9 not 6 sorry umm it is d

OpenStudy (onepieceftw):

multiply the terms and simplify like terms. how can x^s-3x+3x-9 be simplified?

OpenStudy (onepieceftw):

x^2*

OpenStudy (anonymous):

hold on lemme work it out

OpenStudy (onepieceftw):

\[x \times x = x^2\] \[x \times -3 = -3x\] \[3 \times x = 3x\] \[3 \times -3 = -9\]

OpenStudy (onepieceftw):

-3x +3x =0

OpenStudy (anonymous):

yeah thats what the attachmeny says

OpenStudy (onepieceftw):

So x^2-9 is your simplified form

OpenStudy (anonymous):

co it is d

OpenStudy (anonymous):

so*

OpenStudy (anonymous):

i sqaid that earlier

OpenStudy (onepieceftw):

Yes. This question can also be written as (x-3)^2 which makes it a bit easier. I don't know if you've learned that yet.

OpenStudy (onepieceftw):

I wanted to you to do the work without guessing.

OpenStudy (anonymous):

i wasnt guessing i actualy thpught it was d

OpenStudy (onepieceftw):

wanted you*